OpenStudy (anonymous):

which ordered pair is the solution to the system of equations below? x+3y=7 x+2y=10 a.7/2, 13/4 b.7/2, 17/5 c.-2,3 d.16,-3

OpenStudy (alexwee123):

substitute x+3y=7 into x+2y=10 by isolating the x term in x+3y=7 to find x

OpenStudy (alexwee123):

then after you find x plug that x value into an equation to find the y value

OpenStudy (anonymous):

d

OpenStudy (anonymous):